### 24 Jan 2024 v1

# Prompt Design and Engineering: Introduction and Advanced Methods

Abstract
Prompt design and engineering has become an important discipline in just the past few months. In this paper, we provide an introduction to the main concepts as well as review basic and more advanced approaches to prompt design and engineering.
1 Introduction
1.1 What is a prompt?
A prompt in generative AI models is the textual input provided by users to guide the model’s output. This could range from simple questions to detailed descriptions or specific tasks. In the context of image generation models like DALLE-2, prompts are often descriptive, while in LLMs like GPT-3, they can vary from simple queries to complex problem statements.
Prompts generally consist of instructions, questions, input data, and examples. In practice, to elicit a desired response from an AI model, a prompt must contain either instructions or questions, with other elements being optional.
Basic prompts in LLMs can be as simple as asking a direct question or providing instructions for a specific task. Advanced prompts involve more complex structures, such as "chain of thought" prompting, where the model is guided to follow a logical reasoning process to arrive at an answer.
1.2 Basic prompt examples
In order to obtain a result, either 1 or 2 must be present. Everything else is optional. Let’s see a few examples (all of them using ChatGPT).
1.2.1 Instructions + Question
Beyond asking a simple question, possibly the next level of sophistication in a prompt is to include some instructions on how the model should answer the question. Here I ask for advice on how to write a college essay, but also include instructions on the different aspects I am interested to hear about in the answer.
“How should I write my college admission essay? Give me suggestions about the different sections I should include, what tone I should use, and what expressions I should avoid.”
See output in Figure 1
1.2.2 Instructions + Input
Continuing with the previous example, it does not escape anyone that if you can ask a LLM for advice on how to write an essay, you can also directly ask it to write the essay itself.111Note that I am not advocating for this to be an ethical use of these tools, but it is important to be aware that this possibility exists and is already being used by students around the world. It is beyond the scope of this introductory guide to discuss all the possible ethical, legal, or moral concerns that LLMs or generative AI as a whole introduces, but I thought it would be important to at least call out in an introductory example. The fact that you CAN do something with a generative model does not mean that it is the right thing to do! On the other hand, if you are on the receiving end, you better prepare yourself and your organization for all kinds of AI-generated content to come your way. Fortunately, for situations like the one outlined in this example, there are already efforts underway to detect AI generated content.
Let’s see what happens when I input some data about me and give some instructions:
“Given the following information about me, write a 4 paragraph college essay: I am originally from Barcelona, Spain. While my childhood had different traumatic events, such as the death of my father when I was only 6, I still think I had quite a happy childhood.. During my childhood, I changed schools very often, and attended all kinds of schools, from public schools to very religious private ones. One of the most “exotic” things I did during those years is to spend a full school year studying 6th grade in Twin Falls, Idaho, with my extended family.
I started working very early on. My first job, as an English teacher, was at age 13. After that, and throughout my studies, I worked as a teacher, waiter, and even construction worker.”
See output in Figure 2
1.2.3 Question + Examples
You can also feed examples into a language model. In the example below I include some of the shows I like and don’t like to build a “cheap” recommender system. Note that while I added only a few shows, the length of this list is only limited by whatever token limit we might have in the LLM interface.
“Here are some examples of TV shows I really like: Breaking Bad, Peaky Blinders, The Bear. I did not like Ted Lasso. What other shows do you think I might like?”
See output in Figure 3
1.3 Prompt Engineering
Prompt engineering in generative AI models is a rapidly emerging discipline that shapes the interactions and outputs of these models. At its core, a prompt is the textual interface through which users communicate their desires to the model, be it a description for image generation in models like DALLE-2, or a complex problem statement in Large Language Models (LLMs) like GPT-3 and ChatGPT. The prompt can range from simple questions to intricate tasks, encompassing instructions, questions, input data, and examples to guide the AI’s response.
The essence of prompt engineering lies in crafting the optimal prompt to achieve a specific goal with a generative model. This process is not only about instructing the model but also involves a deep understanding of the model’s capabilities and limitations, and the context within which it operates. In image generation models, for instance, a prompt might be a detailed description of the desired image, while in LLMs, it could be a complex query embedding various types of data.
Prompt engineering transcends the mere construction of prompts; it requires a blend of domain knowledge, understanding of the AI model, and a methodical approach to tailor prompts for different contexts. This might involve creating templates that can be programmatically modified based on a given dataset or context. For example, generating personalized responses based on user data might use a template that is dynamically filled with relevant information.
Furthermore, prompt engineering is an iterative and exploratory process, akin to traditional software engineering practices such as version control and regression testing. The rapid growth of this field suggests its potential to revolutionize certain aspects of machine learning, moving beyond traditional methods like feature or architecture engineering, especially in the context of large neural networks. On the other hand, traditional engineering practices such as version control and regression testing need to be adapted to this new paradigm just like they were adapted to other machine learning approaches [1].
This paper aims to delve into this burgeoning field, exploring both its foundational aspects and its advanced applications.
2 Advanced Techniques in Prompt Engineering
2.1 Chain of Thought (CoT)
The Chain of Thought (CoT) technique, initially described in the paper “Chain-of-Thought Prompting Elicits Reasoning in Large Language Models”[2] by Google researchers, represents a pivotal advancement in prompt engineering for Large Language Models (LLMs). This approach hinges on the understanding that LLMs, while proficient in token prediction, are not inherently designed for explicit reasoning. CoT addresses this by guiding the model through essential reasoning steps.
CoT is based on making the implicit reasoning process of LLMs explicit. By outlining the steps required for reasoning, the model is directed closer to a logical and reasoned output, especially in scenarios demanding more than simple information retrieval or pattern recognition.
CoT prompting manifests in two primary forms:
1.
Zero-Shot CoT: This form involves instructing the LLM to “think step by step”, prompting it to deconstruct the problem and articulate each stage of reasoning.
2.
Manual CoT: A more complex variant, it requires providing step-by-step reasoning examples as templates for the model. While yielding more effective results, it poses challenges in scalability and maintenance.
Manual CoT is more effective than zero-shot. However, the effectiveness of this example-based CoT depends on the choice of diverse examples, and constructing prompts with such examples of step by step reasoning by hand is hard and error prone. That is where automatic CoT[3] comes into play.
2.2 Tree of Thought (ToT)
The Tree of Thought (ToT)[4] prompting technique is inspired by the concept of considering various alternative solutions or thought processes before converging on the most plausible one. ToT is based on the idea of branching out into multiple "thought trees" where each branch represents a different line of reasoning. This method allows the LLM to explore various possibilities and hypotheses, much like human cognitive processes where multiple scenarios are considered before determining the most likely one.
A critical aspect of ToT is the evaluation of these reasoning paths. As the LLM generates different branches of thought, each is assessed for its validity and relevance to the query. This process involves real-time analysis and comparison of the branches, leading to a selection of the most coherent and logical outcome.
ToT is particularly useful in complex problem-solving scenarios where a single line of reasoning might not suffice. It allows LLMs to mimic a more human-like problem-solving approach, considering a range of possibilities before arriving at a conclusion. This technique enhances the model’s ability to handle ambiguity, complexity, and nuanced tasks, making it a valuable tool in advanced AI applications.
2.3 Tools, Connectors, and Skills
In the realm of advanced prompt engineering, the integration of Tools, Connectors, and Skills significantly enhances the capabilities of Large Language Models (LLMs). These elements enable LLMs to interact with external data sources and perform specific tasks beyond their inherent capabilities, greatly expanding their functionality and application scope.
Tools in this context are external functions or services that LLMs can utilize. These tools extend the range of tasks an LLM can perform, from basic information retrieval to complex interactions with external databases or APIs.
Connectors act as interfaces between LLMs and external tools or services. They manage data exchange and communication, enabling effective utilization of external resources. The complexity of connectors can vary, accommodating a wide range of external interactions.
Skills refer to specialized functions that an LLM can execute. These encapsulated capabilities, such as text summarization or language translation, enhance the LLM’s ability to process and respond to prompts, even without direct access to external tools.
In the paper “Toolformer: Language Models Can Teach Themselves to Use Tools”[5], the authors go beyond simple tool usage by training an LLM to decide what tool to use when, and even what parameters the API needs. Tools include two different search engines, or a calculator. In the following examples, the LLM decides to call an external Q&A tool, a calculator, and a Wikipedia Search Engine More recently, researchers at Berkeley have trained a new LLM called Gorilla[6] that beats GPT-4 at the use of APIs, a specific but quite general tool.
2.4 Automatic Multi-step Reasoning and Tool-use (ART)
Automatic Multi-step Reasoning and Tool-use (ART)[7] is a prompt engineering technique that combines automated chain of thought prompting with the use of external tools. ART represents a convergence of multiple prompt engineering strategies, enhancing the ability of Large Language Models (LLMs) to handle complex tasks that require both reasoning and interaction with external data sources or tools.
ART involves a systematic approach where, given a task and input, the system first identifies similar tasks from a task library. These tasks are then used as examples in the prompt, guiding the LLM on how to approach and execute the current task. This method is particularly effective when tasks require a combination of internal reasoning and external data processing or retrieval.
2.5 Self-Consistency
Self-Consistency[8] utilizes an ensemble-based method, where the LLM is prompted to generate multiple responses to the same query. The consistency among these responses serves as an indicator of their accuracy and reliability.
The Self-Consistency approach is grounded in the principle that if an LLM generates multiple, similar responses to the same prompt, it is more likely that the response is accurate. This method involves asking the LLM to tackle a query multiple times, each time analyzing the response for consistency. This technique is especially useful in scenarios where factual accuracy and precision are paramount.
The consistency of responses can be measured using various methods. One common approach is to analyze the overlap in the content of the responses. Other methods may include comparing the semantic similarity of responses or employing more sophisticated techniques like BERT-scores or n-gram overlaps. These measures help in quantifying the level of agreement among the responses generated by the LLM.
Self-Consistency has significant applications in fields where the veracity of information is critical. It is particularly relevant in scenarios like fact-checking, where ensuring the accuracy of information provided by AI models is essential. By employing this technique, prompt engineers can enhance the trustworthiness of LLMs, making them more reliable for tasks that require high levels of factual accuracy.
2.6 Reflection
Reflection[9] involves prompting LLMs to assess and potentially revise their own outputs based on reasoning about the correctness and coherence of their responses. The concept of Reflection centers on the ability of LLMs to engage in a form of self-evaluation. After generating an initial response, the model is prompted to reflect on its own output, considering factors like factual accuracy, logical consistency, and relevance. This introspective process can lead to the generation of revised or improved responses.
A key aspect of Reflection is the LLM’s capacity for self-editing. By evaluating its initial response, the model can identify potential errors or areas of improvement. This iterative process of generation, reflection, and revision enables the LLM to refine its output, enhancing the overall quality and reliability of its responses.
2.7 Expert Prompting
Expert Prompting[10] enhances the capabilities of Large Language Models (LLMs) by simulating the responses of experts in various fields. This method involves prompting the LLMs to assume the role of an expert and respond accordingly, providing high-quality, informed answers. A key strategy within Expert Prompting is the multi-expert approach. The LLM is prompted to consider responses from multiple expert perspectives, which are then synthesized to form a comprehensive and well-rounded answer. This technique not only enhances the depth of the response but also incorporates a range of viewpoints, reflecting a more holistic understanding of the subject matter.
2.8 Chains
Chains refer to the method of linking multiple components in a sequence to handle complex tasks with Large Language Models (LLMs). This approach involves creating a series of interconnected steps or processes, each contributing to the final outcome. The concept of Chains is based on the idea of constructing a workflow where different stages or components are sequentially arranged. Each component in a Chain performs a specific function, and the output of one serves as the input for the next. This end-to-end arrangement allows for more complex and nuanced processing, as each stage can be tailored to handle a specific aspect of the task. Chains can vary in complexity and structure, depending on the requirements. In “PromptChainer: Chaining Large Language Model Prompts through Visual Programming”[11], the authors not only describe the main challenges in designing chains, but also describe a visual tool to support those tasks.
2.9 Rails
Rails in advanced prompt engineering refer to a method of guiding and controlling the output of Large Language Models (LLMs) through predefined rules or templates. This approach is designed to ensure that the model’s responses adhere to certain standards or criteria, enhancing the relevance, safety, and accuracy of the output. The concept of Rails involves setting up a framework or a set of guidelines that the LLM must follow while generating responses. These guidelines are typically defined using a modeling language or templates known as Canonical Forms, which standardize the way natural language sentences are structured and delivered.
Rails can be designed for various purposes, depending on the specific needs of the application:
Topical Rails: Ensure that the LLM sticks to a particular topic or domain.
Fact-Checking Rails: Aimed at minimizing the generation of false or misleading information.
Jailbreaking Rails: Prevent the LLM from generating responses that attempt to bypass its own operational constraints or guidelines.
2.10 Automatic Prompt Engineering (APE)
Automatic Prompt Engineering (APE)[12] focuses on automating the process of prompt creation for Large Language Models (LLMs). APE seeks to streamline and optimize the prompt design process, leveraging the capabilities of LLMs themselves to generate and evaluate prompts. APE involves using LLMs in a self-referential manner where the model is employed to generate, score, and refine prompts. This recursive use of LLMs enables the creation of high-quality prompts that are more likely to elicit the desired response or outcome.
The methodology of APE can be broken down into several key steps:
Prompt Generation: The LLM generates a range of potential prompts based on a given task or objective.
Prompt Scoring: Each generated prompt is then evaluated for its effectiveness, often using criteria like clarity, specificity, and likelihood of eliciting the desired response.
Refinement and Iteration: Based on these evaluations, prompts can be refined and iterated upon, further enhancing their quality and effectiveness.
3 Augmenting LLMs through external knowledge - RAG
One of the main limitations of pre-trained LLMs is their lack of up-to-date knowledge or access to private or use-case-specific information. This is where retrieval augmented generation (RAG) comes into the picture [13]. RAG, illustrated in figure 12, involves extracting a query from the input prompt and using that query to retrieve relevant information from an external knowledge source (e.g. a search engine or a knowledge graph, see figure 13 ). The relevant information is then added to the original prompt and fed to the LLM in order for the model to generate the final response.
3.1 RAG-aware prompting techniques
Because of the importance of RAG to build advanced LLM systems, several RAG-aware prompting techniques have been developed recently. One such technique is Forward-looking Active Retrieval Augmented Generation (FLARE)
Forward-looking Active Retrieval Augmented Generation (FLARE)[16] enhances the capabilities of Large Language Models (LLMs) by iteratively combining prediction and information retrieval. FLARE represents an evolution in the use of retrieval-augmented generation, aimed at improving the accuracy and relevance of LLM responses.
FLARE involves an iterative process where the LLM actively predicts upcoming content and uses these predictions as queries to retrieve relevant information. This method contrasts with traditional retrieval-augmented models that typically retrieve information once and then proceed with generation. In FLARE, this process is dynamic and ongoing throughout the generation phase. In FLARE, each sentence or segment generated by the LLM is evaluated for confidence. If the confidence level is below a certain threshold, the model uses the generated content as a query to retrieve relevant information, which is then used to regenerate or refine the sentence. This iterative process ensures that each part of the response is informed by the most relevant and current information available.
For more details on RAG and its relevant works, we refer the readers to this survey of retrieval augmented generations [17].
4 LLM Agents
The idea of AI agents has been well-explored in the history of AI. An agent is typically an autonomous entity that can perceive the environment using its sensors, make a judgment based on the state it currently is, and accordingly act based on the actions that are available to it.
In the context of LLMs, an agent refers to a system based on a specialized instantiation of an (augmented) LLM that is capable of performing specific tasks autonomously. These agents are designed to interact with users and environment to make decisions based on the input and the intended goal of the interaction. Agents are based on LLMs equipped with the ability to access and use tools, and to make decisions based on the given input. They are designed to handle tasks that require a degree of autonomy and decision-making, typically beyond simple response generation.
The functionalities of an LLM-based agent involves include:
Tool Access and Utilization: Agents have the capability to access external tools and services, and to utilize these resources effectively to accomplish tasks.
Decision Making: They can make decisions based on the input, context, and the tools available to them, often employing complex reasoning processes.
As an example, an LLM that has access to a function (or an API) such as weather API, can answer any question related to the weather of the specific place. In other words, it can use APIs to solve problems. Furthermore, if that LLM has access to an API that allows to make purchases, a purchasing agent can be built to not only have capabilities to read information from the external world, but also act on it [18].
4.1 Prompt engineering techniques for agents
Like RAG and Tools, prompt engineering techniques that specifically address the needs of LLM-based agents have been developed. Three such examples are Reasoning without Observation (ReWOO), Reason and Act (ReAct), and Dialog-Enabled Resolving Agents (DERA).
4.1.1 Reasoning without Observation (ReWOO)
Reasoning without Observation (ReWOO)[19] aims to decouple reasoning from direct observations. ReWOO operates by enabling LLMs to formulate comprehensive reasoning plans or meta-plans without immediate reliance on external data or tools. This approach allows the model to create a structured framework for reasoning that can be executed once the necessary data or observations are available.
In ReWOO, the LLM initially develops a plan or a series of steps that outline how to approach and solve a given problem. This meta-planning phase is crucial as it sets the stage for how the model will process information once it becomes available. The execution phase then involves integrating actual data or observations into the pre-formulated plan, leading to a more coherent and contextually relevant response. ReWOO offers significant advantages in terms of token efficiency and robustness to tool failure. It enables LLMs to handle tasks where immediate access to external data is not available, relying instead on a well-structured reasoning framework. This method is particularly advantageous in scenarios where data retrieval is costly, slow, or uncertain, allowing the LLM to maintain a high level of performance and reliability.
4.1.2 Reason and Act (ReAct)
Reason and Act (ReAct)[20] prompts LLMs to generate not only verbal reasoning but also actionable steps, enhancing the model’s dynamic problem-solving capabilities. ReAct is grounded in the principle of integrating reasoning with action. In this approach, the LLM is prompted to alternate between generating reasoning traces (explanations) and actions (steps or commands) in an interleaved manner. This approach allows the model to dynamically reason about a problem and propose concrete actions simultaneously.
4.2 Dialog-Enabled Resolving Agents (DERA)
Dialog-Enabled Resolving Agents (DERA)[21] involves the creation of specialized AI agents that can engage in dialogue, resolve queries, and make decisions based on interactive exchanges. DERA is structured around the idea of utilizing multiple agents within a dialog context, each with specific roles and functions. These agents can include Researchers, who gather and analyze information, and Deciders, who make final judgments based on the information provided. This division of roles allows for a more organized and efficient approach to problem-solving and decision-making.
DERA is particularly advantageous in scenarios requiring complex decision-making and problem-solving, such as in medical diagnostics or customer service. The collaborative and interactive nature of DERA agents allows them to handle intricate queries with a level of depth and nuance that single-agent systems might struggle with. Moreover, this approach aligns closely with human decision-making processes, making the AI’s reasoning and conclusions more relatable and trustworthy.
5 Prompt Engineering Tools and Frameworks
The development of advanced prompt engineering techniques has been accompanied by the emergence of various tools and frameworks designed to facilitate the implementation and optimization of these methods. Below is an overview of some of the prominent tools and frameworks in this domain.
Langchain222 is a widely recognized prompt engineering toolkit. Initially focused on supporting Chains, it has evolved to support Agents and various Tools for diverse functionalities, ranging from memory handling to web browsing.
Developed by Microsoft, Semantic Kernel333 is a toolkit designed for skills and planning in C# and Python. It now encompasses capabilities beyond its initial scope, including chaining, indexing, memory access, and plugin development.
Guidance444, another library from Microsoft, utilizes a templating language to support a variety of prompt engineering techniques. Being a more recent addition to the field, it offers modern solutions and supports the methods discussed in this guide.
Developed by NVidia, Nemo Guardrails555 is a tool aimed at building Rails to ensure LLMs operate within desired parameters and guidelines.
LlamaIndex666 focuses on managing the data that feeds into LLM applications, primarily through data connectors and tools.
From Intel, FastRAG777 includes not only the basic Retrieval Augmented Generation (RAG) approach but also more advanced implementations of RAG, aligning with techniques discussed in this guide.
Auto-GPT888 is a popular tool for designing LLM agents, facilitating the creation and management of sophisticated AI agents. More recently, textbfAutoGen by Microsoft has gained popularity for designing agents and multi-agent systems.
In summary, these tools and frameworks play a crucial role in the practical application and advancement of prompt engineering, offering solutions that range from basic prompt management to complex AI agent design and operation.
6 Conclusion
In just a few months prompt design and engineering has become an extremely important aspect of designing LLM and Generative AI systems. In this paper, we have introduced the basics as well as more advanced methods. The space is rapidly evolving and we should expect to see a lot of innovation to happen very quickly. This introduction and guide should still serve as a good guide with an early historical perspective.